# Performance of Quantum Approximate Optimization with Quantum Error Detection

### v1

@@ V.3 Approximation ratio and logical fidelity
To relate the logical fidelity and the noisy approximation ratio we consider that the circuit noise takes the form of a global white noise channel [62]
@@
Moreover, we sometimes use the white noise channel to estimate the logical fidelity from the samples obtained by running the unencoded circuits and the Iceberg code circuits on hardware or the emulator. In this scenario we have access to the experimental approximation ratio, or equivalently, to the average energies $\braket{\mathcal{H}}_{c}$ for every QAOA circuit $c\in C_{\rm une}\cup C_{\rm ice}$ considered in this work, unencoded or protected by the Iceberg code. The estimated logical fidelity is the one that reproduces such average energy under the white noise channel:
@@
When we have access to the experimental implementation of the QAOA circuits on hardware or on the emulator this is the estimated logical fidelity we report in this work.
@@
obtained from the experimental expected values $\braket{Z_{i}Z_{j}}_{c}$. Note that in general these experimental ratios are not expected to take the same value for all edges, but under the white noise channel, they all equal the channel fidelity.
To quantify the deviation between the experimental data and the white noise simplification, we consider the normalized distance between the experimental logical fidelity and the ratios as
@@ V.4 Model fitting
For fitting the Iceberg code model to the experimental data, still leveraging the least squares method, we minimize the residuals of both the logical fidelity and discard rate. Given a QAOA instance $c\in C_{\rm ice}$ with experimental discard rate $D_{c}$ we define the loss function for multiple QAOA instances as
@@ Acknowledgments
We would like to thank the experimental scientists in the Quantinuum team that design and maintain the H-series devices. We extend our gratitude to Drs. Matthew DeCross, Ciaran Ryan-Anderson and Selwyn Simsek for valuable discussions about the modeling of memory errors in these devices and Drs. David Hayes, Changhao Li, Pradeep Niroula, and Shree Hari Sureshbabu for the preliminary statistical analysis of the numerical experiments. We appreciate the comments and suggestions of Drs. Chris N. Self, Michael Perlin and Sivaprasad Omanakuttan on the manuscript. We thank the technical staff at JPMorganChase’s Global Technology Applied Research Center for their support and helpful discussions.
@@ Appendix A Iceberg QAOA circuits details
In this appendix we depict the error detection blocks of the Iceberg code and briefly describe the strategy taken from [25] to mitigate memory noise.
@@
Error detection blocks of the Iceberg code. The initialization in Fig. 7, taken from [63], prepares all logical qubits in the initial logical state |+¯$\ket{\bar{+}}^{\otimes k}$. If any of the two ancilla measurements (indicated in red) is $-1$, an error is detected and the circuit is discarded. The syndrome measurement in Fig. 8, and the final measurement in Fig. 9 were taken from [18]. Both circuits measure the two Iceberg code stabilizers in order to detect errors, indicated again by excited ancilla measurements. The measurement of the $S_{Z}$ stabilizer is performed in the post-process of the $n$ destructive measurement outcomes of the code qubits, where its value is the parity of the outputs. If no errors are detected, the logical outcomes are extracted as well from pairwise parities. These circuits are fault-tolerant, so that every single error from their components (cnot, single-qubit gates, qubit initialization, qubit measurement, or idle qubits) either acts trivially without causing a logical error, or is detectable. Consequently, at least two errors are necessary to cause a logical error.
@@ Appendix B Experimental details
Gap between hardware and emulator. In Table 4, we compare the performance of H2-1 device and emulator experiments on both Iceberg and unencoded circuits. These QAOA circuits are fixed with $\ell=10$ and have varying number $k$ of logical qubits. The Iceberg circuits use four syndrome measurements. The differences in logical fidelity between H2-1 device and emulator are small for unencoded circuits but become increasingly significant for Iceberg code circuits as $k$ increases, where the hardware moderately outperforms the emulator. The emulator consistently estimates larger values of the post-selection ratio. The H2-1 hardware experiments in Fig. 1C show that for up to $k=20$ and $\ell=10$ with four syndrome measurements, the Iceberg circuits outperforms the unencoded circuits. However, this performance is beyond the frontier observed in the H2-1 emulator experiments and the model predictions. This discrepancy is likely due to the fidelity gap between the H2-1 device and emulator when handling relatively large-sized circuits that have both large qubit count and circuit depth.
@@
Appendix C Model details
[…]
Justifying Simplification 2. This simplification can be numerically tested for small number $k$ of logical qubits and number $\mu$ of faulty gates. Table 5 shows a comparison of the simplified model fractions against these numerical values. For the initialization block, the fractions can be read directly from Table 2. For the syndrome and the final measurement blocks only harmless $H$ input errors are considered, so the model fractions can be read directly from the first columns of Tables 6 and 7. The table shows small deviations between the model and the numerical fractions, with the model tending to overestimate the fraction of undetectable errors ($H$ and $L$). However, the deviations reduce quickly as the number $k$ of logical qubits or the number $\mu$ of faulty gates increases.
Syndrome measurement block. Table 6 summarises the contributions from input and internal errors. For undetectable input errors internal errors cause 16 possible excitation events depending on which of the two ancillas and two stabilizers are excited or not. Since input exciting errors produce 3 possible ancilla excitations, the number of combined events grows to 48. If the ancillas excited by the input exciting error and by internal errors coincide, the circuit is not discarded at this block.
Final measurement block. Contributions are summarized in Table 7. For input undetectable errors internal errors produce 8 excitation events depending on which of the two stabilizers and the flag are excited or not. For input exciting errors the number of excitation events grows to 24. Finally, since no errors can be propagated to the next block every detectable error at this block produces an immediate discard. We consequently set the update rule $E\leftarrow 0$. From the contributions in top of the table we obtain the update rules in the bottom part of the table.
@@ Appendix C Model details
Table 5: Comparison the fractions of errors that contribute to the four block probabilities, as predicted by the model and as calculated exactly for small number of logical qubits and faulty gates. Only the first significant digits are reported but the values are exact; they do not carry any uncertainty.

### v3

@@ V.3 Approximation ratio and logical fidelity
To relate the logical fidelity and the noisy approximation ratio, we consider that the circuit noise takes the form of a global white noise channel [69]
@@
Moreover, we sometimes use the white noise channel to estimate the logical fidelity from the samples obtained by running the unencoded circuits and the Iceberg code circuits on hardware or the emulator. In this scenario, we have access to the experimental approximation ratio, or equivalently, to the average energies $\braket{\mathcal{H}}_{c}$ for every QAOA circuit $c\in C_{\rm une}\cup C_{\rm ice}$ considered in this work, unencoded or protected by the Iceberg code. The estimated logical fidelity is the one that reproduces such average energy under the white noise channel:
@@
When we have access to the experimental implementation of the QAOA circuits on hardware or the emulator, this is the estimated logical fidelity we report in this work.
@@
obtained from the experimental expected values $\braket{Z_{i}Z_{j}}_{c}$. Note that in general, these experimental ratios are not expected to take the same value for all edges, but under the white noise channel, they all equal the channel fidelity.
To quantify the deviation between the emulated data and the white noise simplification, we consider the normalized distance between the experimental logical fidelity and the ratios as
@@ V.4 Model fitting
Still leveraging the least squares method to fit the Iceberg code model to the emulated data, we minimize the residuals of both the logical fidelity and discard rate. Given a QAOA instance $c\in C_{\rm ice}$ with experimental discard rate $D_{c}$, we define the loss function for multiple QAOA instances as
@@
Data availability
Data for reproducing all figures and tables is available at is available at https://zenodo.org/records/15231805.
Code availability
The code for numerical simulations is available upon reasonable request.
@@ Acknowledgments
We would like to thank the experimental scientists in the Quantinuum team that design and maintain the H-series devices. We extend our gratitude to Drs. Matthew DeCross, Ciaran Ryan-Anderson and Selwyn Simsek for valuable discussions about the modeling of memory errors in these devices and Drs. David Hayes, Changhao Li, Pradeep Niroula, and Shree Hari Sureshbabu for the preliminary statistical analysis of the numerical experiments. We appreciate the comments and suggestions of Drs. Chris N. Self, Michael Perlin and Sivaprasad Omanakuttan on the manuscript. We thank the technical staff at JPMorganChase’s Global Technology Applied Research Center for their support and helpful discussions. This paper was prepared for informational purposes with contributions from the Global Technology Applied Research center of JPMorgan Chase & Co. This paper is not a product of the Research Department of JPMorgan Chase & Co. or its affiliates. Neither JPMorgan Chase & Co. nor any of its affiliates makes any explicit or implied representation or warranty and none of them accept any liability in connection with this paper, including, without limitation, with respect to the completeness, accuracy, or reliability of the information contained herein and the potential legal, compliance, tax, or accounting effects thereof. This document is not intended as investment research or investment advice, or as a recommendation, offer, or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction.
Author Contributions
Z.H. developed emulation and model fitting pipeline. D.A. developed the performance model and conducted the hardware experiments. R.S. contributed to the data analysis and experimental design. All authors participated in technical discussions and contributed to the writing of the manuscript.
Competing Interests
The authors declare no competing financial or non-financial interests.
@@ Appendix A Iceberg QAOA circuits details
In this appendix, we depict the error detection blocks of the Iceberg code and briefly describe the strategy taken from [26] to mitigate memory noise.
@@
Error detection blocks of the Iceberg code. The initialization in Fig. 7, taken from [70], prepares all logical qubits in the initial logical state |+¯$\ket{\bar{+}}^{\otimes k}$. If any of the two ancilla measurements (indicated in red) is $-1$, an error is detected and the circuit is discarded. The syndrome measurement in Fig. 8 and the final measurement in Fig. 9 were taken from [19]. Both circuits measure the two Iceberg code stabilizers in order to detect errors, indicated again by excited ancilla measurements. The measurement of the $S_{Z}$ stabilizer is performed in the post-process of the $n$ destructive measurement outcomes of the code qubits, where its value is the parity of the outputs. If no errors are detected, the logical outcomes are extracted as well from pairwise parities. These circuits are fault-tolerant, so that every single error from their components (cnot, single-qubit gates, qubit initialization, qubit measurement, or idle qubits) either acts trivially without causing a logical error, or is detectable. Consequently, at least two errors are necessary to cause a logical error.
@@ Appendix B Experimental details
Gap between hardware and emulator. In Table 4, we compare the performance of H2-1 device and emulator experiments on both Iceberg and unencoded circuits. These QAOA circuits are fixed with $\ell=10$ and have varying number $k$ of logical qubits. The Iceberg circuits use four syndrome measurements. The differences in logical fidelity between H2-1 device and emulator are small for unencoded circuits but become increasingly significant for Iceberg code circuits as $k$ increases, where the hardware moderately outperforms the emulator. The emulator consistently estimates larger values of the post-selection ratio. The H2-1 hardware experiments in Fig. 1C show that for up to $k=20$ and $\ell=10$ with four syndrome measurements, the Iceberg circuits outperform the unencoded circuits. However, this performance is beyond the frontier observed in the H2-1 emulator experiments and the model predictions. This discrepancy is likely due to the fidelity gap between the H2-1 device and emulator when handling relatively large-sized circuits that have both large qubit count and circuit depth.
@@
Runtime Estimation. It is also intriguing to examine the runtime of the Iceberg QAOA circuit compared to an unencoded circuit. Since samples may be discarded during the execution of the Iceberg program if fault-tolerant gadgets detect errors, let $t_{ps}$ represent the execution time of a circuit where all sample discarding is deferred to the end of the program. We introduce a factor $\lambda$ to quantify the runtime reduction of a program with discarding, denoted as $t_{ice}$: {align} t_ice & = λ* t_ps
λ = ∑_i^g d_i t_i / t_ps where $g$ is the total number of FT gadets (including FT initialization, syndrome measurement, and FT final measurement), $d_{i}$ is the ratio of samples that end execution at the $i$-th fault-tolerant (FT) gadget with $\sum_{i}^{g}d_{i}=1$, and $t_{i}$ is the runtime up to the $i$-th gadget. As the last gadget is the FT final measurement, we have $t_{g}=t_{ps}$.
While the actual hardware execution time of programs is typically not accessible to users, we estimate the execution time of the circuits using data from the H2-1 emulator. This emulator reports the duration of a single circuit shot by accounting for both quantum gate operations and classical operations, such as ion movement and cooling. Additionally, by tracking all syndrome measurements during the emulations, we can readily calculate the ratios $d_{i}$.
@@
For the five instances listed in Table. 4 with $k=$16, 18, 20, 22, 24, their runtime results are summarized in Table. 6. The non-monotonic change in the runtime of unencoded circuits $t_{une}$ as $k$ increases is attributed to the sub-optimality of the pytket compilation pass. For these instances, the time per shot of an encoded circuit, $t_{ice}$, is approximately 1.9 to 3 times greater than $t_{une}$. Additionally, more total shots are required to obtain a sufficient number of post-selected shots. For example, in the emulation experiments of these five instances, we set a total of 3000 shots for encoded circuits and 500 shots for unencoded ones in emulations, resulting in an overall slowdown of 12 to 18 times for Iceberg QAOA compared to the unencoded version. This demonstrates a trade-off between computational time and improved performance.
While the hardware execution runtime of a QED/QEC encoded circuit is a critical consideration for practical deployment, it is important to note that we do not currently claim any speedup for QED circuits. The primary purpose of applying Iceberg encoding is to aid in benchmarking the QAOA algorithm on hardware. We have provided a basic estimation of execution time above, but a more comprehensive study on runtime is deferred to future work.
Perturbation level $\delta$ 0 0.1 0.2 0.3 0.5 1
Noiseless 0.988 0.969 0.924 0.855 0.659 0.544
Encoded H2-1E 0.958 $\pm$ 0.005 0.955 $\pm$ 0.004 0.905 $\pm$ 0.005 0.832 $\pm$ 0.006 0.659 $\pm$ 0.006 0.542 $\pm$ 0.005
Unencoded H2-1E 0.943 $\pm$ 0.003 0.930 $\pm$ 0.004 0.890 $\pm$ 0.006 0.833 $\pm$ 0.006 0.649 $\pm$ 0.006 0.543 $\pm$ 0.005
Table 5: Performance of $\ell=10$ QAOA with perturbed parameters on a $k=16$ graph instance: The values following the $\pm$ symbol represent standard errors.
Appendix C Additional experiments
[…]
Figure 12: Emulated performance of Iceberg QAOA and unencoded QAOA on an SK problem with $k=10$. Iceberg QAOA, implemented with $s=4$, outperforms the unencoded one when $\ell\geq 3$, as accumulated noise starts to negatively impact the unencoded QAOA.
Iceberg QAOA on the Sherrington-Kirkpatrick problem. This paper primarily demonstrates the performance of Iceberg QAOA on the MaxCut problem, but the approach can be readily extended to other problem families. In addition, we showcase the application of Iceberg QAOA on the Sherrington-Kirkpatrick (SK) model, another widely used benchmark for quantum optimization algorithms. The problem Hamiltonian is given by:
$\mathcal{H}=\sum_{i<j}J_{i,j}Z_{i}Z_{j},$ (24)
where $J_{i,j}$ follows a standard normal distribution. We employ both Iceberg and unencoded QAOA to solve a $k=10$ SK problem instance using a set of fixed parameters as reported in [74].
As illustrated in Fig. 12, when varying the depth $\ell$, Iceberg QAOA with $s=4$ significantly outperforms the unencoded version. At $\ell=1$, their performance is similar, as the circuit is too shallow to manifest the effects of noise.
Appendix D Model details
[…]
Justifying Simplification 2. This simplification can be numerically tested for small number $k$ of logical qubits and number $\mu$ of faulty gates. Table 7 shows a comparison of the simplified model fractions against these numerical values. For the initialization block, the fractions can be read directly from Table 2. For the syndrome and the final measurement blocks, only harmless $H$ input errors are considered, so the model fractions can be read directly from the first columns of Tables 8 and 9. The table shows small deviations between the model and the numerical fractions, with the model tending to overestimate the fraction of undetectable errors ($H$ and $L$). However, the deviations reduce quickly as the number $k$ of logical qubits or the number $\mu$ of faulty gates increases.
Syndrome measurement block. Table 8 summarises the contributions from input and internal errors. For undetectable input errors, internal errors cause 16 possible excitation events depending on which of the two ancillas and two stabilizers are excited or not. Since input exciting errors produce 3 possible ancilla excitations, the number of combined events grows to 48. If the ancillas excited by the input exciting error and by internal errors coincide, the circuit is not discarded at this block.
Final measurement block. Contributions are summarized in Table 9. For input undetectable errors, internal errors produce 8 excitation events depending on which of the two stabilizers and the flag are excited or not. For input exciting errors, the number of excitation events grows to 24. Finally, since no errors can be propagated to the next block, every detectable error at this block produces an immediate discard. We consequently set the update rule $E\leftarrow 0$. From the contributions at the top of the table, we obtain the update rules in the bottom part of the table.
$k$ 16 18 20 22 24
$\lambda$ 0.692 0.633 0.588 0.523 0.513
$t_{ps}(s)$ 2.6 3.4 3.6 4.8 5.5
$t_{ice}(s)$ 1.8 2.2 2.3 2.5 2.8
$t_{une}(s)$ 0.6 0.9 1.2 1.3 1.1
Table 6: Estimated runtime of the five instances shown in Table. 4.
@@ Appendix D Model details
Table 7: A comparison of the fractions of errors that contribute to the four block probabilities, as predicted by the model and as calculated exactly for a small number of logical qubits and faulty gates. Only the first significant digits are reported but the values are exact; they do not carry any uncertainty.